What is the SAMPLE STANDARD DEVIATION (use your calculator) of this data set?

4, 6, 11, 6, 8, 7

(Round to the 2nd decimal place: Example 7.32
1
Expert's answer
2020-10-26T19:54:09-0400

The sample size is  "n=6." The provided sample data along with the data required to compute the sample mean "\bar{x}" and sample variance "s^2" are shown in the table below:


"\begin{matrix}
 & x & x^2 \\
 & 4 & 16 \\
 & 6 & 36 \\
 & 6 & 36 \\
 & 7 & 49 \\
 & 8 & 4 \\
 & 11 & 121 \\
Sum= & 42 & 322
\end{matrix}"

The sample mean "\bar{x}" is computed as follows:


"\bar{x}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nx_i=\dfrac{42}{6}=7"

The sample variance "s^2" is


"s^2=\dfrac{1}{n-1}\bigg(\displaystyle\sum_{i=1}^nx_i^2-\dfrac{1}{n}\big(\displaystyle\sum_{i=1}^nx_i\big)^2\bigg)="

"=\dfrac{1}{6-1}\big(322-\dfrac{1}{6}(42)^2\big)=5.6"

Therefore, the sample sandartd deviation "s" is


"s=\sqrt{s^2}=\sqrt{5.6}\approx2.37"